# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ IV.2 Nonlocality of the noisy TBSM distributions
[…]
V Discussion
We have derived noise-robust proofs of quantum nonlocality in the triangle network without inputs. Our key ingredient is a result of approximate rigidity for parity token (PTC) local distributions. This allowed us to obtain bounds on the admissible noise for a general family of quantum distributions that we introduced, generalizing the RGB4 distributions [14]. We characterized noise-robustness in terms of limited visibility for the source states, considering dephasing and white noise. Alternatively, we can also characterize distributions in the vicinity of a target distribution, via a bound on the total-variation distance.
@@ V Discussion
Another direction is to see how to strengthen the noise robustness for the nonlocality proof in our Result 4. A promising option is to revisit Result 3, and strengthen the last condition (3.iii). We used the standard implementation of the inflation technique, but stronger versions could be obtained by exploiting nonlinear constraints. While we could prove robustness to white noise up to $0.67\%$, there is still potential for significant improvements; indeed numerical methods [17] suggests robustness of up to $\sim 10\%$. Any progress in this direction would be relevant, in particular from the perspective of experimental demonstrations. Concerning dephasing noise, we demonstrated strong robustness, up to $\sim 80\%$. We conjecture that there exist quantum distributions that can tolerate an amount of dephasing noise arbitrarily close to one.
It would also be interesting to see if the property of approximate PTC rigidity can be extended from local models to quantum ones. If not, this could indicate the existence of quantum nonlocal correlations in the triangle network with binary outputs (and no inputs).
Code availability.— Our implementation of the method for the triangle network can be found at https://github.com/SadraBoreiri/Robust-Nework-Nonlocality
@@ Appendix A Approximate PTC rigidity
Consider variables $\alpha$, $\beta$,$\gamma$ sampled from the sets $A,B,C$ with some probability distributions/densities, and deterministic response function $a(\beta,\gamma),b(\alpha,\gamma),c(\alpha,\beta)\in\{0,1\}$. Let there be a subset $\Lambda\subset A\times B\times C$ such that
@@
$G_{\gamma}=\{(\alpha,\beta)|(\alpha,\beta,\gamma)\in\Lambda\}\subset A\times B$ (49)
@@
as well as the value $T_{\gamma_{*}}=1$. It is easy to see that on the set $G_{*}\times\{\gamma_{*}\}\subset\Lambda$ these function satisfy the PTC rigidity conditions given by
@@
Remaining in the slice $\gamma$, can define the following binary functions
@@
For both choices $z=0$ and $z=1$, there functions satisfy the PTC conditions on the set $S_{\gamma}\times\{\gamma\}$
@@
$A_{\gamma}^{z}$ =$=\{\alpha|T_{\alpha}\oplus t_{\alpha}^{z}=0\}$ (65)
$B_{\gamma}^{z}$ =$=\{\alpha|T_{\beta}\oplus t_{\beta}^{z}=0\},$ (66)
such that $A_{\gamma}^{0}$ with $A_{\gamma}^{1}$ and $B_{\gamma}^{0}$ with $B_{\gamma}^{1}$ are disjoint, and $S_{\gamma}\subset(A_{\gamma}^{0}\cup A_{\gamma}^{1})\times(B_{\gamma}^{0}\cup B_{\gamma}^{1})$. In fact for any point $(\alpha,\beta)\in S_{\gamma}$ it is true that
@@
because it is both in $G_{*}$ and $G_{\gamma}$. Hence, any such point is either in $A_{\gamma}^{0}\times B_{\gamma}^{0}$ or $A_{\gamma}^{1}\times B_{\gamma}^{1}$, and we conclude that
@@
This allows us to split $S_{\gamma}=S_{\gamma}^{0}\cup S_{\gamma}^{1}$ in two disjoint sets with $S_{\gamma}^{z}\subset A_{\gamma}^{z}\times B_{\gamma}^{z}$. Not that for the choice $t_{\gamma}=z$ the token functions $T_{\alpha}$ and $T_{\gamma}$ are PTC on $S_{\gamma}^{z}\times\{\gamma\}$. And our goal is of course to pick the value of $z$ for which $\text{Pr}(S_{\gamma}^{z})$ is maximized. Let us define this probability
@@
We will now lower-bound this probability. First, not that there is a trivial bound given by $\hat{P}_{\gamma}\geq\frac{1-\delta_{\gamma}}{2}$. Furthermore, we have
@@
To this this note that for $z^{\prime}=\text{argmin}_{z}\text{Pr}(S^{z}_{\gamma})$ we have $\text{Pr}(S_{\gamma}^{z^{\prime}})\leq\text{Pr}(A_{\gamma}^{z^{\prime}}\times B_{\gamma}^{z^{\prime}})$, and $\text{Pr}(S_{\gamma}^{z^{\prime}})\leq\text{Pr}(S_{\gamma}^{z^{\prime}\oplus 1})\leq\text{Pr}(A_{\gamma}^{z^{\prime}\oplus 1}\times B_{\gamma}^{z^{\prime}\oplus 1})$ (since $S^{z}_{\gamma}\subset A^{z}_{\gamma}\times B_{\gamma}^{z}$). In addition, we know that
$1-\delta_{\gamma}=P_{\gamma}=\text{Pr}(S_{\gamma}^{0})+\text{Pr}(S_{\gamma}^{1})\leq\text{Pr}(A_{\gamma}^{0}\times B_{\gamma}^{0})+\text{Pr}(A_{\gamma}^{1}\times B_{\gamma}^{1}).$ (71)
Denoting $p_{z}=\text{Pr}(A_{\gamma}^{z})$ and $q_{z}=\text{Pr}(A_{\gamma}^{z})$ with $p_{0}+p_{1},q_{0}+q_{1}\leq 1$, we want to upper bound $\min_{z}\text{Pr}(A^{z}_{\gamma}\times B_{\gamma}^{z})$ subject to the constraint in Eq. (71). This is equivalent to solving the following min-max problem
@@
In the section A.2 we solve this optimization, proving the following bound
$\min_{z}\text{Pr}(A_{\gamma}^{z}\times B_{\gamma}^{z})\leq\left(\frac{1-\sqrt{1-2\delta_{\gamma}}}{2}\right)^{2}\quad\implies\quad\hat{P}_{\gamma}\geq 1-\delta_{\gamma}-\left(\frac{1-\sqrt{1-2\delta_{\gamma}}}{2}\right)^{2}.$ (73)
@@
on which we have shown a consistent assignment of the token functions $T_{\alpha},T_{\gamma},T_{\gamma}$ fulfilling the PTC condition (for $\gamma=\gamma_{*}$ we, of course, have $\hat{S}_{\gamma^{*}}=G_{*}$).
@@
[…]
@@
Figure 7: The function $f(\delta_{\gamma})$.
@@
by $\varepsilon_{*}\leq\varepsilon$. In the next section we derive a tighter upper-bound which uses the fact that $\delta_{\gamma}\geq\varepsilon_{*}$.
@@ A.1 Tighter bound on $\text{Pr}(\Upsilon)$
A tight convex lower bound on $f(\delta_{\gamma})$ depicted in Fig. 7 on the interval $[\varepsilon_{*},1]$ is given by the piece-wise linear function which connects connecting the points $\{\varepsilon_{*},f_{*}=f(\varepsilon_{*})\}$ with $\{1/2,1/4\}$ for $\delta_{\gamma}\in[\varepsilon_{*},1/2]$ and the points $\{1/2,1/4\}$ with $\{1,0\}$ for $\delta_{\gamma}\in[1/2,1]$. Formally, this lower bound is given by
@@
and $\text{Pr}(\Upsilon)\geq f^{\varepsilon_{*}}_{\text{con}}(\varepsilon+\varepsilon_{*})$ for any scenario with $\text{Pr}(G_{*})=1-\varepsilon_{*}$. But since $\varepsilon_{*}$ is not known we need to take the worst-case lower bound
@@
illustrated in Fig. 8. As a final remark consider the first order expansion $1-\text{f}_{\varepsilon}=1-\frac{5}{2}\varepsilon+O(\varepsilon^{2})$ to see that for very small $\varepsilon$ the two lower bounds are quite different.
@@
$\begin{split}\max_{p_{0},p_{1},q_{0},q_{1}}&\quad p_{0}q_{0}\\ \text{such that}&\quad p_{0}q_{0}\leq p_{1}q_{1}\\ &\quad p_{0}q_{0}\geq 1-\delta_{\gamma}-p_{1}q_{1}\\ &\quad p_{0}\leq 1-p_{1}\\ &\quad q_{0}\leq 1-q_{1}\\ &\quad p_{0},p_{1},q_{0},q_{1}\geq 0.\\ \end{split}$ (97)
The variables $p_{1}$ and $q_{1}$ now do not appear in the goal function. Furthermore, the constraints on $p_{0}$ and $q_{0}$ are easier to satisfy, the larger values of $p_{1}$, $q_{1}$ and $p_{1}q_{1}$. Hence the maximum is attained when these auxiliary variables take the maximal values $p_{1}=1-p_{0}$ and $q_{1}=1-p_{0}$. This allows us to rewrite the maximization in terms of two variables $p=p_{0}$ and $q=q_{0}$ as follows
$\begin{split}\max_{p,q\in[0,1]}&\quad p\,q\\ \text{such that}&\quad p\,q\leq(1-p)(1-q)\\ &\quad p\,q\geq 1-\delta_{\gamma}-(1-p)(1-q)\end{split}\qquad\text{or}\qquad\begin{split}\max_{p,q\in[0,1]}&\quad p\,q\\ \text{such that}&\quad p+q\leq 1\\ &\quad 2\,p\,q\geq p+q-\delta_{\gamma}\end{split}.$ (98)
Let us introduce the variables $g=\sqrt{p\,q}$ and $a=\frac{p+q}{2}$ which are the geometric and the algebraic means of $p,q$. We there help we write
$\begin{split}\max_{g,a}&\quad g^{2}\\ \text{such that}&\quad a\leq\frac{1}{2}\\ &\quad g^{2}\geq a-\frac{1}{2}\delta_{\gamma},\end{split}$ (99)
and we also know that the algebraic and geometric means satisfy $a\geq g$, which can be saturated $a=g$ by the choice $p=q$. Note that for a fixed free variable $a$, the choice $g=a$ is the one that maximizes the goal function and relaxes the constraints the most. We can thus set $a=g$ without loss of generality in order to obtain
@@ Appendix B The noisy W-distribution
and the other terms are left unspecified. It satisfies the PTC condition with probability $1-\varepsilon$. Hence, by our Result 2 we know that there must exist a local PTC distribution $P(a,b,c)$ such that $\delta(P_{W}^{(\varepsilon,\delta)},P)\leq\text{f}_{\varepsilon}$. With straightforward algebra, this condition can rewritten as
@@
since there must exist a combination of $q_{\alpha},q_{\beta},q_{\gamma}$ for which $\Delta^{(\varepsilon,\delta)}(P)\leq 0$ holds. To prove that a distribution $P_{W}^{(\varepsilon,\delta)}$ is not triangle local it is thus sufficient to show that $\Delta_{min}^{(\varepsilon,\delta)}>0$.
@@
To perform the minimization in Eq. (106), we used the Nelder-Mead or simplex search method [30] via the Python library SciPy [31]. The results of the minimization are reported in Fig. 9. Note that this is an empirical minimization algorithm, which does not provide a convergence guarantee. Therefore, we have only proven the nonlocality of a given distribution under the assumption that the global minimum was found. Nevertheless, the goal function $\Delta^{(\varepsilon,\delta)}(q_{\alpha},q_{\beta},q_{\gamma})$ is a smooth and well-behaved function of three variables on a compact interval, and with more computational resources it is possible to do the minimization in a guaranteed manner, e.g. using a branch-and-bound algorithm.

### v3

@@ 5.2 Nonlocality of the noisy TBSM distributions
Naturally, this can be combined with Result 3 in order to show that no local distribution $\widetilde{P}$ fulfilling (5.i)(5.iii)(5.i)-(5.iii) can exist. The argument for this is fairly similar to the corollary of Result 4. In Appendix F we present an LP relaxation of this feasibility problem. In Fig. 8 we plot the size of the ball $\varepsilon$ around the TBSM distributions with $\lambda_{0}^{2}=0.22$ which we prove to be nonlocal. The largest value we find is $\varepsilon=0.24\%$ and it is achieved for an RGB4 distribution ($w=0$).
6 Discussion
We have derived noise-robust proofs of quantum nonlocality in the triangle network without inputs. Our key ingredient is a result of approximate rigidity for parity token (PTC) local distributions. These methods can be applied to an arbitrary noisy distribution. In particular, we illustrate the relevance of the method by obtaining bounds on the admissible noise for a general family of quantum distributions that we introduced, generalizing the RGB4 distributions [16]. We characterized noise-robustness in terms of limited visibility for the source states, considering dephasing and white noise. Alternatively, we can also characterize distributions in the vicinity of a target distribution, via a bound on the total-variation distance.
@@ 6 Discussion
Another direction is to see how to strengthen the noise robustness for the nonlocality proof in our Result 4. A promising option is to revisit Result 3, and strengthen the last condition (3.iii). We used the standard implementation of the inflation technique, but stronger versions could be obtained by exploiting nonlinear constraints. While we could prove robustness to white noise up to $0.55\%$, there is still potential for significant improvements; indeed numerical methods [21] suggest robustness of up to $\sim 10\%$. Any progress in this direction would be relevant, in particular from the perspective of experimental demonstrations. Concerning dephasing noise, we demonstrated strong robustness, up to $\sim 80\%$. We conjecture that there exist quantum distributions that can tolerate any amount of dephasing noise arbitrarily close to one.
It would also be interesting to see if the property of approximate PTC rigidity can be extended from local models to quantum ones. If not, one can hope to find a quantum distribution violating the approximate PTC rigidity, indicating the existence of quantum nonlocal correlations within the triangle network with binary outputs and no inputs.
Code availability.— Our implementation of the method for the triangle network can be found at https://github.com/SadraBoreiri/Robust-Network-Nonlocality
@@ Appendix A Approximate PTC rigidity
In a triangle-local model, consider variables $\alpha$, $\beta$, $\gamma$ sampled from the sets $\mathit{A,B,\Gamma}$ respectively, according to some probability distributions. Without loss of generality, we can assume that the model includes deterministic response functions $a(\beta,\gamma),b(\alpha,\gamma),c(\alpha,\beta)\in\{0,1\}$ [^6] For any indeterministic local response function at the parties we can simply move the indeterminacy to the sources $\alpha$, $\beta$, $\gamma$..
Approximate PTC rigidity states that, for any triangle-local model leading to a binary output distribution $P(a,b,c)$ satisfying the approximate parity condition $\text{Pr}(a\oplus b\oplus c=1)\geq 1-\varepsilon$, there exists a subset $\Upsilon\subset\mathit{A}\times\mathit{B}\times\mathit{\Gamma}$ of local variables with high probability $Pr(\bm{\xi}\in\Upsilon)\geq 1-O(\varepsilon)$, within which it is feasible to define the token functions $T_{\alpha},T_{\beta},T_{\gamma}$ satisfying the PTC rigidity $a(\beta,\gamma)=T_{\beta}\oplus T_{\gamma}\oplus 1$ etc. We prove this here, in two main steps. In the first step, we characterize the subset $\Upsilon$ and show the consistent assignment of the token functions inside this subset. In the second step, we prove that this subset happens with high probability, specifically $Pr(\Upsilon)\geq 1-3\varepsilon$. Moreover, in Section A.1 we prove a slightly more involved and tighter lower bound, showing that $Pr(\Upsilon)\geq 1-\text{f}_{\varepsilon}=1-\frac{5}{2}\varepsilon+O(\varepsilon^{2})$.
Step 1: Defining a subset $\Upsilon$ of local variables with consistent assignment of token functions
Here our objective is to specify $\Upsilon\subset\mathit{A}\times\mathit{B}\times\mathit{\Gamma}$ a subset of local variables on which we can consistently assign the token functions $T_{\alpha},T_{\beta},T_{\gamma}$ satisfying the PTC condition
To do so, first let $\Lambda$ be the subset of $\mathit{A}\times\mathit{B}\times\mathit{\Gamma}$ in which the outputs satisfy the PTC condition, i.e.
@@
$G_{\gamma}=\{(\alpha,\beta)|(\alpha,\beta,\gamma)\in\Lambda\}\subset\mathit{A}\times\mathit{B}$ (45)
@@
as well as the value $T_{\gamma_{*}}=1$. It is easy to see that on the set $G_{*}\times\{\gamma_{*}\}\subset\Lambda$ these functions satisfy the PTC rigidity conditions given by
@@
Remaining in the slice $\gamma$, we can define the following binary functions
@@
For both choices $z=0$ and $z=1$, these functions satisfy the PTC conditions on the set $S_{\gamma}\times\{\gamma\}$
@@
$\mathit{A}_{\gamma}^{z}$ $=\{\alpha|T_{\alpha}\oplus t_{\alpha}^{z}=0\}$ (61)
$\mathit{B}_{\gamma}^{z}$ $=\{\alpha|T_{\beta}\oplus t_{\beta}^{z}=0\},$ (62)
such that $\mathit{A}_{\gamma}^{0}$ with $\mathit{A}_{\gamma}^{1}$ and $\mathit{B}_{\gamma}^{0}$ with $\mathit{B}_{\gamma}^{1}$ are disjoint, and $S_{\gamma}\subset(\mathit{A}_{\gamma}^{0}\cup\mathit{A}_{\gamma}^{1})\times(\mathit{B}_{\gamma}^{0}\cup\mathit{B}_{\gamma}^{1})$. In fact for any point $(\alpha,\beta)\in S_{\gamma}$ it is true that
@@
because it is both in $G_{*}$ and $G_{\gamma}$. Hence, any such point is either in $\mathit{A}_{\gamma}^{0}\times\mathit{B}_{\gamma}^{0}$ or $\mathit{A}_{\gamma}^{1}\times\mathit{B}_{\gamma}^{1}$, and we conclude that
@@
This allows us to split $S_{\gamma}=S_{\gamma}^{0}\cup S_{\gamma}^{1}$ in two disjoint sets with $S_{\gamma}^{z}\subset\mathit{A}_{\gamma}^{z}\times\mathit{B}_{\gamma}^{z}$. Note that for the choice $t_{\gamma}=z$ the token functions $T_{\alpha}$ and $T_{\gamma}$ are PTC on $S_{\gamma}^{z}\times\{\gamma\}$. Our goal is of course to pick the value of $z$ for which $\text{Pr}(S_{\gamma}^{z})$ is maximized. Let us define this probability
@@
We will now lower-bound this probability. First, note that there is a trivial bound given by $\hat{P}_{\gamma}\geq\frac{1-\delta_{\gamma}}{2}$. Furthermore, we have
@@
To this note that for $z^{\prime}=\text{argmin}_{z}\text{Pr}(S^{z}_{\gamma})$ we have $\text{Pr}(S_{\gamma}^{z^{\prime}})\leq\text{Pr}(\mathit{A}_{\gamma}^{z^{\prime}}\times\mathit{B}_{\gamma}^{z^{\prime}})$, and $\text{Pr}(S_{\gamma}^{z^{\prime}})\leq\text{Pr}(S_{\gamma}^{z^{\prime}\oplus 1})\leq\text{Pr}(\mathit{A}_{\gamma}^{z^{\prime}\oplus 1}\times\mathit{B}_{\gamma}^{z^{\prime}\oplus 1})$ (since $S^{z}_{\gamma}\subset\mathit{A}^{z}_{\gamma}\times\mathit{B}_{\gamma}^{z}$). In addition, we know that
$1-\delta_{\gamma}=P_{\gamma}=\text{Pr}(S_{\gamma}^{0})+\text{Pr}(S_{\gamma}^{1})\leq\text{Pr}(\mathit{A}_{\gamma}^{0}\times\mathit{B}_{\gamma}^{0})+\text{Pr}(\mathit{A}_{\gamma}^{1}\times\mathit{B}_{\gamma}^{1}).$ (67)
Denoting $p_{z}=\text{Pr}(\mathit{A}_{\gamma}^{z})$ and $q_{z}=\text{Pr}(\mathit{A}_{\gamma}^{z})$ with $p_{0}+p_{1},q_{0}+q_{1}\leq 1$, we want to upper bound $\min_{z}\text{Pr}(\mathit{A}^{z}_{\gamma}\times\mathit{B}_{\gamma}^{z})$ subject to the constraint in Eq. (67). This is equivalent to solving the following min-max problem
@@
In Section A.2 we solve this optimization, proving the following bound
$\min_{z}\text{Pr}(\mathit{A}_{\gamma}^{z}\times\mathit{B}_{\gamma}^{z})\leq\left(\frac{1-\sqrt{1-2\delta_{\gamma}}}{2}\right)^{2}\quad\implies\quad\hat{P}_{\gamma}\geq 1-\delta_{\gamma}-\left(\frac{1-\sqrt{1-2\delta_{\gamma}}}{2}\right)^{2}.$ (69)
@@
on which we have shown a consistent assignment of the token functions $T_{\alpha},T_{\beta},T_{\gamma}$ fulfilling the PTC condition (for $\gamma=\gamma_{*}$ we, of course, have $\hat{S}_{\gamma^{*}}=G_{*}$).
Step 2: proving $\text{Pr}(\Upsilon)\geq 1-3\varepsilon$.
@@
$\text{Pr}(\Omega)=1-\text{Pr}(\bm{\xi}\notin\Lambda\text{ or }(\alpha,\beta)\notin G_{*})\geq 1-\text{Pr}(\bm{\xi}\notin\Lambda)-\text{Pr}((\alpha,\beta)\notin G_{*})\geq 1-\varepsilon-\varepsilon_{*}.$ (74)
@@
Figure 9: The function $f(\delta_{\gamma})$ evaluated in the interval $[0,1]$. The concavity of the function can be observed in the $[0,1/2]$ interval.
@@
by $\varepsilon_{*}\leq\varepsilon$. In the next Section, we derive a tighter upper bound which uses the fact that $\delta_{\gamma}\geq\varepsilon_{*}$.
@@ A.1 Tighter bound on $\text{Pr}(\Upsilon)$
A tight convex lower bound on $f(\delta_{\gamma})$ depicted in Fig. 9 on the interval $\varepsilon_{*},1]$ is given by the piecewise linear function which connects the points $\{\varepsilon_{*},f_{*}=f(\varepsilon_{*})\}$ with $\{1/2,1/4\}$ for $\delta_{\gamma}\in[\varepsilon_{*},1/2]$ and the points $\{1/2,1/4\}$ with $\{1,0\}$ for $\delta_{\gamma}\in[1/2,1]$. Formally, this lower bound is given by
@@
and $\text{Pr}(\Upsilon)\geq f^{\varepsilon_{*}}_{\text{con}}(\varepsilon+\varepsilon_{*})$ for any scenario with $\text{Pr}(G_{*})=1-\varepsilon_{*}$. But since $\varepsilon_{*}$ is not known, we need to take the worst-case lower bound
@@
illustrated in Fig. 10. As a final remark, consider the first order expansion $1-\text{f}_{\varepsilon}=1-\frac{5}{2}\varepsilon+O(\varepsilon^{2})$ to see that for very small $\varepsilon$ the two lower bounds are quite different.
@@
[…]
The variables $p_{1}$ and $q_{1}$ now do not appear in the goal function. Furthermore, the constraints on $p_{0}$ and $q_{0}$ are easier to satisfy with larger values of $p_{1}$, $q_{1}$ and $p_{1}q_{1}$. Hence the maximum is attained when these auxiliary variables take the maximal values $p_{1}=1-p_{0}$ and $q_{1}=1-p_{0}$. This allows us to rewrite the maximization in terms of two variables $p=p_{0}$ and $q=q_{0}$ as follows
$\begin{split}\max_{p,q\in[0,1]}&\quad p\,q\\ \text{such that}&\quad p\,q\leq(1-p)(1-q)\\ &\quad p\,q\geq 1-\delta_{\gamma}-(1-p)(1-q)\end{split}\qquad\text{which can be rewritten as}\qquad\begin{split}\max_{p,q\in[0,1]}&\quad p\,q\\ \text{such that}&\quad p+q\leq 1\\ &\quad 2\,p\,q\geq p+q-\delta_{\gamma}\end{split}.$ (94)
Let us introduce the variables $g=\sqrt{p\,q}$ and $a=\frac{p+q}{2}$ which are the geometric and the algebraic means of $p,q$. With their help, we write the following optimization and they are equivalent as knowing $a,g$ gives $p,q$ as well with $p,q=a\pm\sqrt{a^{2}-g^{2}}$
$\begin{split}\max_{g,a}&\quad g^{2}\\ \text{such that}&\quad a\leq\frac{1}{2}\\ &\quad g^{2}\geq a-\frac{1}{2}\delta_{\gamma},\\ &\quad a\geq g.\end{split}$ (95)
The final constraint arises because the algebraic and geometric means satisfy the inequality $a\geq g$, and guarantees that $p$ and $q$ are well defined, which can be saturated $a=g$ by the choice $p=q$. Note that for a fixed free variable $a$, the choice $g=a$ is the one that maximizes the goal function and relaxes the constraints the most. We can thus set $a=g$ without loss of generality in order to obtain
@@ Appendix B The noisy W-distribution
[…]
@@
since there must exist a combination of $q_{\alpha},q_{\beta},q_{\gamma}$ for which $\Delta^{(\varepsilon,\delta)}(P)\leq 0$ holds. To prove that a distribution $P_{W}^{(\varepsilon,\delta)}$ is not triangle-local it is thus sufficient to show that $\Delta_{min}^{(\varepsilon,\delta)}>0$.
@@
To perform the minimization in Eq. (102), we used the Nelder-Mead or simplex search method [38] via the Python library SciPy [39]. The results of the minimization are reported in Fig. 11. Note that this is an empirical minimization algorithm, which does not provide a convergence guarantee. Note that the goal function $\Delta^{(\varepsilon,\delta)}(q_{\alpha},q_{\beta},q_{\gamma})$ is a smooth and well-behaved function of three variables on a compact interval; therefore, it is plausible that the solver has found its global minimum. Nevertheless, it should be stressed that this only proves the nonlocality of a given distribution under the assumption that the global minimum was indeed found. In the future, it is desirable to perform the same minimization with more computational resources and the guarantee of convergence. This can e.g. be done with the help of the branch-and-bound algorithm, given that the goal function is Lipschitz-continuous on a compact three-parameter subset.